# Graph Limits via Quotients

### v1

@@ 3.3 Duality between Grapheurs and Graphons
In this section, we characterize convergence of grapheurs in terms of graphons and vice-versa, exhibiting a facet of the duality between the two graph limit theories. Since graphon theory is most simply expressed in the undirected case (see [diaconis2007graph, §8] for the directed case), we restrict ourselves to this setting in the present section.
We begin by recalling a few basic notions from graphon theory. Let $\widetilde{\mathcal{W}}=\{W\colon[0,1]^{2}\to[0,1]\textrm{ measurable},\allowbreak\textrm{ symmetric}\}$ be the space of $[0,1]$-valued kernels. We define two kernels $W$ and $U$ to be equivalent, denoted $W\sim U$, if there exist (Lebesgue-)measure-preserving maps $\varphi,\psi\colon[0,1]\to[0,1]$ satisfying $W\circ(\varphi,\varphi)=U\circ(\psi,\psi)$. This is the weak isomorphism between kernels commonly defined in the dense graph limits literature [lovasz2012large, Cor. 10.35]. We denote by $\mathcal{W}=\widetilde{\mathcal{W}}$$\sim$ the quotient space under this equivalence relation. Elements of $\mathcal{W}$ are called graphons [lovasz2012large, §7.1]. A graphon $W\in\mathcal{W}$ defines random simple graphs of each size by setting $\mathbb{G}_{k}[W]=(\mathsf{B}_{T_{i},T_{j}}^{(W)})_{i,j\in[k]}$ where $\mathsf{B}_{T_{i},T_{j}}^{(W)}\sim\mathrm{Ber}(W(T_{i},T_{j}))$ are drawn independently for $i<j$ while $\mathsf{B}_{T_{j},T_{i}}^{(W)}=\mathsf{B}_{T_{i},T_{j}}^{(W)}$ for $i>j$ and $\mathsf{B}_{T_{i},T_{i}}^{(W)}=0$. Here $T_{1},\ldots,T_{k}\overset{iid}{\sim}\mathrm{Unif}([0,1])$ as before. A sequence $(W_{n})\subseteq\mathcal{W}$ converges if for each $k$, the sequence of random graphs $(\mathbb{G}_{k}[W_{n}])_{k}$ converges weakly for each kk. [^4] Graphon convergence is usually defined by total variation convergence of these random graphs, which is equivalent to their weak convergence since they are supported on the finite set $\mathbb{S}^{k}_{\mathrm{sim}}$. This is equivalent to convergence of homomorphism densities and to convergence in cut metric [lovasz2012large, §10]. Finally, a finite graph $G\in\mathbb{S}^{n}_{\mathrm{sim}}$ defines a step graphon $W_{G}(x,y)=G_{i,j}$ if $x\in I_{i}^{(n)}\times I_{j}^{(n)}$ where $I_{i}^{(n)}=[(i-1)/n,i/n)$.
@@
weakly. Therefore, we get $\langle G,\mathbb{G}_{k}[W_{n}]\rangle\to\langle G,\mathbb{G}_{k}[W]\rangle$ for all $G\in\mathbb{S}^{k}$ by linearity and the fact that $\mathbb{G}_{k}[W_{n}]$ all have zero diagonal. This implies $\mathbb{G}_{k}[W_{n}]\to\mathbb{G}_{k}[W]$ weakly by Cramér–Wold [kallenberg1997foundations, Cor. 6.5]. Since this holds for all $k\in\mathbb{N}$, we conclude that $W_{n}\to W$. Conversely, suppose $W_{n}\to W$. Then $\langle\mathsf{M}_{G},W_{n}\rangle=\langle G,\mathsf{G}_{k}[W_{n}]\rangle\to\langle G,\mathsf{G}_{k}[W]\rangle=\langle\mathsf{M}_{G},W\rangle$ for any $G\in\mathbb{S}^{k}\cap\Delta^{k\times k}$. ∎
@@
Theorem 3.10 exhibits a concrete, formal duality between quotient convergence and graphon convergence. This duality can be extended to general grapheurs and more general kernels on $0,1]^{2}$, but the extension of graphon convergence to this setting is more complicated [diaconis2007graph, §8]. We therefore leave this and other extensions of this duality for future work.
@@ 3.4 Missing Proofs from Section 3
For the first claim, since $\mathsf{M}$ is exchangeable we conclude that $\mu=\mathbb{E}\mathsf{M}$ is a (deterministic) exchangeable measure on $0,1]^{2}$, meaning that $\mu\circ(\sigma,\sigma)=\mu$ for all (Lebesgue)-measure preserving bijections $\sigma\colon[0,1]\to[0,1]$. The space of such measures is two-dimensional and is spanned by $\lambda^{2},\lambda_{D}$ by [kallenberg2005probabilistic, Thm. 9.12], so $\mu=\theta\lambda+\vartheta\lambda_{D}$ for some $\theta,\vartheta\in\mathbb{R}$. Since $\mathsf{M}$ is a probability measure almost surely, its mean $\mu$ is a probability measure as well, necessitating $\theta+\vartheta=1$ and $\theta,\vartheta\geq 0$. This proves the first claim.
@@
We apply the disintegration theorem [disintegration, Thm. A] to the continuous map $\pi_{k}\colon\mathcal{P}([0,1]^{2})\to\Delta^{k\times k}$ between compact metric spaces sending $\mu\mapsto(\mu(I_{i}^{(k)}\times I_{j}^{(k)}))_{i,j\in[k]}$. The disintegration theorem yields for each $\mathsf{M}\in\mathcal{M}$ a Borel collection of probability measures $(\lambda_{G}^{(\mathsf{M})})_{G\in\Delta^{k\times k}}$ for each $\mathsf{M}\in\mathcal{M}$ such that $\mathrm{supp}(\lambda_{G}^{(\mathsf{M})})\subseteq\pi_{k}^{-1}(G)$ and such that $\mathrm{Law}(\mathsf{M})=\mu_{k}^{(\mathsf{M})}\otimes\lambda_{G}^{(\mathsf{M})}$ in the sense that
@@
Next, we turn to proving Proposition 3.3, which states that $\mathcal{M}$ is weakly compact and that the metric $W_{\square}$ metrizes this weak topology. To prove that $\mathcal{M}$ is weakly compact, we rely on the following lemma and the characterization of $\mathcal{M}$ from [kallenberg2005probabilistic, §9] as the space of ergodic random exhangeable measures.
@@
By [kallenberg2005probabilistic, Prop. 9.1], the space of random exchangeable measures is precisely the collection of random measures on $0,1]^{2}$ invariant under the countable group $\widetilde{\mathsf{S}}_{\infty}$ consisting of permutations of the intervals $(I_{i}^{(n)})_{i\in[n]}$ for all $n\in\mathbb{N}$. By [kallenberg2005probabilistic, Lemma A1.2], we can characterize its extreme points by
@@
Endow $\mathbb{R}^{2}$ with the $\ell_{\infty}$ norm, endow $\mathcal{P}([0,1]^{2})$ with the Wasserstein-1 metric with respect to this norm, and endow $\mathcal{M}$ with the Wasserstein-1 metric with respect to the Wasserstein-1 metric on $\mathcal{P}([0,1]^{2})$, which metrizes weak convergence of random measures [villani2008optimal, Thm. 6.9]. Explicitly, this metric is given by
@@
where the min over couplings is attained [villani2008optimal, Thm. 4.1]. It therefore suffices to prove that $\lim_{n}W_{\square}(\mathsf{M}_{n},\mathsf{M})=0$ if and only if $\lim_{n}\mathrm{dist}_{W_{1}}(\mathsf{M}_{n},\mathsf{M})=0$. We proceed to prove this by standard approximations of continuous functions by step functions and vice-versa.
@@ 4.1 Edge Sampling View of Quotient Convergence
obtained by first sampling $\mu_{n}\sim\mathsf{M}_{n}$ and then sampling $(x_{i},y_{i})\overset{iid}{\sim}\mu_{n}$ converge weakly by [kallenberg1973characterization, Thm. 1.1], since the distributions of (15) are precisely the finite-dimensional distributions of $\mathsf{M}_{n}$. The exchangeable array (15) is precisely the list of edges of the multigraph $k\mathsf{G}^{(k)}[\mathsf{M}_{n}]$ listed in arbitrary order after assigning its vertices iid labels in $\mathrm{Unif}([0,1])$. Formally, we have $\mathsf{M}_{\mathsf{G}^{(k)}[\mathsf{M}_{n}]}\overset{d}{=}\frac{1}{k}\sum_{i=1}^{k}\delta_{(x_{i}^{(n)},y_{i}^{(n)})}$ for all $n,k\in\mathbb{N}$. Thus, the sequence (15) converges weakly for each $k\in\mathbb{N}$ precisely when the sequence of random graphs $(\mathsf{G}^{(k)}[\mathsf{M}_{n}])_{n}$ converges weakly. ∎
@@
The random graphs we obtain by sampling edges from a grapheur appeared previously in the literature as edge-exchangeable random multigraph models [janson2018edge]. As explained in [janson2018edge, Rmk. 4.4], every edge-exchangeable random multigraph model is obtained from a random exchangeable measure on $0,1]^{2}$ as above after labelling the vertices by iid labels from $\mathrm{Unif}([0,1])$. Nevertheless, to our knowledge graph limits based on these edge-exchangeable random graphs as in Proposition 4.1 have not been previously studied in the literature, see also Section 1.2.
As noted in [janson2018edge], if we repeat the above edge sampling process infinitely-many times, each edge $(T_{i},T_{j})$ corresponding to a positive edge weight $E_{i,j}>0$ will be repeated infinitely-many times. In contrast, all the other edges sampled above will only appear once. In particular, the vertices $T^{\prime},T_{1}^{\prime},T_{2}^{\prime}$ sampled in this process will only appear in one edge. Such vertices are called “blips” in [borgs2019sampling, janson2018edge], and the isolated edges and self-loops are called “dust”.
@@ 4.2 Edge-Based Szemerédi Regularity
Second, note that $f$ is $1/n$-Lipschitz in each $(x_{i},y_{i})$. Indeed, if $\mathsf{G}^{(n,i)}[\mathsf{M}]$ differs from $\mathsf{G}^{(n)}[\mathsf{M}]$ only in the $i$th location $(x_{i},y_{i})$, and we glue the two couplings $(\mathsf{M}^{\prime\prime},\mathsf{M}^{\prime\prime}_{\mathsf{G}^{(n)}[\mathsf{M}]})$ and $(\mathsf{M}^{\prime\prime},\mathsf{M}^{\prime\prime}_{\mathsf{G}^{(n,i)}[\mathsf{M}]})$ attaining the corresponding $\overline{W_{\square}}$ distances as in [villani2008optimal, §1], then
@@
since the only difference between $\mathsf{M}^{\prime\prime}_{\mathsf{G}^{(n,i)}[\mathsf{M}]}$ and $\mathsf{M}^{\prime\prime}_{\mathsf{G}^{(n)}[\mathsf{M}]}$ is a weight of $1/n$ placed at one of their atoms. Since $(x_{1},y_{1}),\ldots,(x_{n},y_{n})$ are sampled iid from a fixed measure $\mu$ (after it was sampled once from $\mathsf{M}$), the bounded-difference concentration inequality [wainwright2019high, Cor. 2.21] implies that $f(\mathsf{G}^{(n)}[\mathsf{M}])\leq\mathbb{E}f(\mathsf{G}^{(n)}[\mathsf{M}])+\epsilon/\sqrt{n}$ with probability at least $1-\exp(-2\epsilon^{2})$, proving the theorem. ∎
@@
We conclude this section by remarking that the above discussion applies to undirected random graphs sampled from symmetric grapheurs.
@@ 4.3 Property Testing
Suppose $f$ is quotient-testable. Define $\mathsf{M}_{<\infty}=\{\mathsf{M}_{G}:G\in\Delta^{n\times n}\textrm{ for some }n\}\to\mathbb{R}$ to be the grapheurs in $\mathsf{M}$ arising from finite graphs and define $\bar{f}\colon\mathsf{M}_{<\infty}\to\mathbb{R}$ by $\bar{f}(\mathsf{M}_{G})=f(G)$. This is well-defined because $f$ is permutation-invariant and isolate-indifferent, as $\mathsf{M}_{G}=\mathsf{M}_{G^{\prime}}$ if and only if $G$ and $G^{\prime}$ differ by isolated vertices and node relabeling [kallenberg1990exchangeable, Prop. 3]. Further, $\bar{f}$ is continuous by continuity of $f$, hence it extends uniquely to a continuous function on the closure $\mathcal{M}$ of $\mathcal{M}_{<\infty}$. This proves statement 2.
@@
For the first claim, it was shown in [levin2025deFin, Thm. 5.2] that there exists a polynomial $q_{k}$ over $\mathbb{R}^{k\times k}$ satisfying $f(G_{n})=\mathbb{E}q_{k}(\rho(F_{k,n})G_{n})$ for all $G_{n}\in\Delta^{n\times n}$ and all $n\in\mathbb{N}$. Since $q_{k}$ is $L$-Lipschitz on $\Delta^{k\times k}$ for some $L>0$, we have by Proposition 3.2 that
@@
We remark that the graphon analog of the characterization of quotient-testable parameters in Theorem 4.10 was used as the definition of testable parameters in [convergent_seqs1, Def. 2.11].
@@
where $S_{2}^{(\mathrm{out})},S_{2}^{(\mathrm{in})}\in\mathbb{N}^{3\times 3}$ are outward- and inward-pointing stars with two leaves, respectively. Define the polynomial $q_{2}$ on $2\times 2$ matrices by

### v2

@@ 3.3 Duality between Grapheurs and Graphons
In this section, we characterize convergence of grapheurs in terms of graphons and vice-versa, exhibiting a facet of the duality between the two graph limit theories. Since graphon theory is most simply expressed in the undirected case (see [17, §8] for the directed case), we restrict ourselves to this setting in the present section.
We begin by recalling a few basic notions from graphon theory. Let $\widetilde{\mathcal{W}}=\{W\colon[0,1]^{2}\to[0,1]\textrm{ measurable},\allowbreak\textrm{ symmetric}\}$ be the space of $[0,1]$-valued kernels. We define two kernels $W$ and $U$ to be equivalent, denoted $W\sim U$, if there exist (Lebesgue-)measure-preserving maps $\varphi,\psi\colon[0,1]\to[0,1]$ satisfying $W\circ(\varphi,\varphi)=U\circ(\psi,\psi)$. This is the weak isomorphism between kernels commonly defined in the dense graph limits literature [34, Cor. 10.35]. We denote by $\mathcal{W}=\widetilde{\mathcal{W}}$$\sim$ the quotient space under this equivalence relation. Elements of $\mathcal{W}$ are called graphons [34, §7.1]. A graphon $W\in\mathcal{W}$ defines random simple graphs of each size by setting $\mathbb{G}_{k}[W]=(\mathsf{B}_{T_{i},T_{j}}^{(W)})_{i,j\in[k]}$ where $\mathsf{B}_{T_{i},T_{j}}^{(W)}\sim\mathrm{Ber}(W(T_{i},T_{j}))$ are drawn independently for $i<j$ while $\mathsf{B}_{T_{j},T_{i}}^{(W)}=\mathsf{B}_{T_{i},T_{j}}^{(W)}$ for $i>j$ and $\mathsf{B}_{T_{i},T_{i}}^{(W)}=0$. Here $T_{1},\ldots,T_{k}\overset{iid}{\sim}\mathrm{Unif}([0,1])$ as before. A sequence $(W_{n})\subseteq\mathcal{W}$ converges if for each $k$, the sequence of random graphs $(\mathbb{G}_{k}[W_{n}])_{k}$ converges weakly for each kk. [^4] Graphon convergence is usually defined by total variation convergence of these random graphs, which is equivalent to their weak convergence since they are supported on the finite set $\mathbb{S}^{k}_{\mathrm{sim}}$. This is equivalent to convergence of homomorphism densities and to convergence in cut metric [34, §10]. Finally, a finite graph $G\in\mathbb{S}^{n}_{\mathrm{sim}}$ defines a step graphon $W_{G}(x,y)=G_{i,j}$ if $x\in I_{i}^{(n)}\times I_{j}^{(n)}$ where $I_{i}^{(n)}=[(i-1)/n,i/n)$.
@@
weakly. Therefore, we get $\langle G,\mathbb{G}_{k}[W_{n}]\rangle\to\langle G,\mathbb{G}_{k}[W]\rangle$ for all $G\in\mathbb{S}^{k}$ by linearity and the fact that $\mathbb{G}_{k}[W_{n}]$ all have zero diagonal. This implies $\mathbb{G}_{k}[W_{n}]\to\mathbb{G}_{k}[W]$ weakly by Cramér–Wold [29, Cor. 6.5]. Since this holds for all $k\in\mathbb{N}$, we conclude that $W_{n}\to W$. Conversely, suppose $W_{n}\to W$. Then $\langle\mathsf{M}_{G},W_{n}\rangle=\langle G,\mathsf{G}_{k}[W_{n}]\rangle\to\langle G,\mathsf{G}_{k}[W]\rangle=\langle\mathsf{M}_{G},W\rangle$ for any $G\in\mathbb{S}^{k}\cap\Delta^{k\times k}$. ∎
@@
Theorem 3.10 exhibits a concrete, formal duality between quotient convergence and graphon convergence. This duality can be extended to general grapheurs and more general kernels on $0,1]^{2}$, but the extension of graphon convergence to this setting is more complicated [17, §8]. We therefore leave this and other extensions of this duality for future work.
@@ 3.4 Missing Proofs from Section 3
For the first claim, since $\mathsf{M}$ is exchangeable we conclude that $\mu=\mathbb{E}\mathsf{M}$ is a (deterministic) exchangeable measure on $0,1]^{2}$, meaning that $\mu\circ(\sigma,\sigma)=\mu$ for all (Lebesgue)-measure preserving bijections $\sigma\colon[0,1]\to[0,1]$. The space of such measures is two-dimensional and is spanned by $\lambda^{2},\lambda_{D}$ by [26, Thm. 9.12], so $\mu=\theta\lambda+\vartheta\lambda_{D}$ for some $\theta,\vartheta\in\mathbb{R}$. Since $\mathsf{M}$ is a probability measure almost surely, its mean $\mu$ is a probability measure as well, necessitating $\theta+\vartheta=1$ and $\theta,\vartheta\geq 0$. This proves the first claim.
@@
We apply the disintegration theorem [39, Thm. A] to the continuous map $\pi_{k}\colon\mathcal{P}([0,1]^{2})\to\Delta^{k\times k}$ between compact metric spaces sending $\mu\mapsto(\mu(I_{i}^{(k)}\times I_{j}^{(k)}))_{i,j\in[k]}$. The disintegration theorem yields for each $\mathsf{M}\in\mathcal{M}$ a Borel collection of probability measures $(\lambda_{G}^{(\mathsf{M})})_{G\in\Delta^{k\times k}}$ for each $\mathsf{M}\in\mathcal{M}$ such that $\mathrm{supp}(\lambda_{G}^{(\mathsf{M})})\subseteq\pi_{k}^{-1}(G)$ and such that $\mathrm{Law}(\mathsf{M})=\mu_{k}^{(\mathsf{M})}\otimes\lambda_{G}^{(\mathsf{M})}$ in the sense that
@@
Next, we turn to proving Proposition 3.3, which states that $\mathcal{M}$ is weakly compact and that the metric $W_{\square}$ metrizes this weak topology. To prove that $\mathcal{M}$ is weakly compact, we rely on the following lemma and the characterization of $\mathcal{M}$ from [26, §9] as the space of ergodic random exhangeable measures.
@@
By [26, Prop. 9.1], the space of random exchangeable measures is precisely the collection of random measures on $0,1]^{2}$ invariant under the countable group $\widetilde{\mathsf{S}}_{\infty}$ consisting of permutations of the intervals $(I_{i}^{(n)})_{i\in[n]}$ for all $n\in\mathbb{N}$. By [26, Lemma A1.2], we can characterize its extreme points by
@@
Endow $\mathbb{R}^{2}$ with the $\ell_{\infty}$ norm, endow $\mathcal{P}([0,1]^{2})$ with the Wasserstein-1 metric with respect to this norm, and endow $\mathcal{M}$ with the Wasserstein-1 metric with respect to the Wasserstein-1 metric on $\mathcal{P}([0,1]^{2})$, which metrizes weak convergence of random measures [46, Thm. 6.9]. Explicitly, this metric is given by
@@
where the min over couplings is attained [46, Thm. 4.1]. It therefore suffices to prove that $\lim_{n}W_{\square}(\mathsf{M}_{n},\mathsf{M})=0$ if and only if $\lim_{n}\mathrm{dist}_{W_{1}}(\mathsf{M}_{n},\mathsf{M})=0$. We proceed to prove this by standard approximations of continuous functions by step functions and vice-versa.
@@ 4.1 Edge Sampling View of Quotient Convergence
obtained by first sampling $\mu_{n}\sim\mathsf{M}_{n}$ and then sampling $(x_{i},y_{i})\overset{iid}{\sim}\mu_{n}$ converge weakly by [27, Thm. 1.1], since the distributions of (15) are precisely the finite-dimensional distributions of $\mathsf{M}_{n}$. The exchangeable array (15) is precisely the list of edges of the multigraph $k\mathsf{G}^{(k)}[\mathsf{M}_{n}]$ listed in arbitrary order after assigning its vertices iid labels in $\mathrm{Unif}([0,1])$. Formally, we have $\mathsf{M}_{\mathsf{G}^{(k)}[\mathsf{M}_{n}]}\overset{d}{=}\frac{1}{k}\sum_{i=1}^{k}\delta_{(x_{i}^{(n)},y_{i}^{(n)})}$ for all $n,k\in\mathbb{N}$. Thus, the sequence (15) converges weakly for each $k\in\mathbb{N}$ precisely when the sequence of random graphs $(\mathsf{G}^{(k)}[\mathsf{M}_{n}])_{n}$ converges weakly. ∎
@@
The random graphs we obtain by sampling edges from a grapheur appeared previously in the literature as edge-exchangeable random multigraph models [24]. As explained in [24, Rmk. 4.4], every edge-exchangeable random multigraph model is obtained from a random exchangeable measure on $0,1]^{2}$ as above after labelling the vertices by iid labels from $\mathrm{Unif}([0,1])$. Nevertheless, to our knowledge graph limits based on these edge-exchangeable random graphs as in Proposition 4.1 have not been previously studied in the literature, see also Section 1.2.
As noted in [24], if we repeat the above edge sampling process infinitely-many times, each edge $(T_{i},T_{j})$ corresponding to a positive edge weight $E_{i,j}>0$ will be repeated infinitely-many times. In contrast, all the other edges sampled above will only appear once. In particular, the vertices $T^{\prime},T_{1}^{\prime},T_{2}^{\prime}$ sampled in this process will only appear in one edge. Such vertices are called “blips” in [7, 24], and the isolated edges and self-loops are called “dust”.
@@ 4.2 Edge-Based Szemerédi Regularity
Second, note that $f$ is $1/n$-Lipschitz in each $(x_{i},y_{i})$. Indeed, if $\mathsf{G}^{(n,i)}[\mathsf{M}]$ differs from $\mathsf{G}^{(n)}[\mathsf{M}]$ only in the $i$th location $(x_{i},y_{i})$, and we glue the two couplings $(\mathsf{M}^{\prime\prime},\mathsf{M}^{\prime\prime}_{\mathsf{G}^{(n)}[\mathsf{M}]})$ and $(\mathsf{M}^{\prime\prime},\mathsf{M}^{\prime\prime}_{\mathsf{G}^{(n,i)}[\mathsf{M}]})$ attaining the corresponding $\overline{W_{\square}}$ distances as in [46, §1], then
@@
since the only difference between $\mathsf{M}^{\prime\prime}_{\mathsf{G}^{(n,i)}[\mathsf{M}]}$ and $\mathsf{M}^{\prime\prime}_{\mathsf{G}^{(n)}[\mathsf{M}]}$ is a weight of $1/n$ placed at one of their atoms. Since $(x_{1},y_{1}),\ldots,(x_{n},y_{n})$ are sampled iid from a fixed measure $\mu$ (after it was sampled once from $\mathsf{M}$), the bounded-difference concentration inequality [47, Cor. 2.21] implies that $f(\mathsf{G}^{(n)}[\mathsf{M}])\leq\mathbb{E}f(\mathsf{G}^{(n)}[\mathsf{M}])+\epsilon/\sqrt{n}$ with probability at least $1-\exp(-2\epsilon^{2})$, proving the theorem. ∎
@@
We remark that $\overline{W_{\square}}$ also metrizes weak converge in $\mathcal{M}$, as can be shown by combining Theorem 4.3 with Proposition 4.1. We opted to use the $W_{\square}$ metric in the rest of the paper since it is easier to relate to random quotients as in Proposition 3.2. We conclude this section by remarking that the above discussion applies to undirected random graphs sampled from symmetric grapheurs.
@@ 4.3 Property Testing
Suppose $f$ is quotient-testable. Define $\mathsf{M}_{<\infty}=\{\mathsf{M}_{G}:G\in\Delta^{n\times n}\textrm{ for some }n\}\to\mathbb{R}$ to be the grapheurs in $\mathsf{M}$ arising from finite graphs and define $\bar{f}\colon\mathsf{M}_{<\infty}\to\mathbb{R}$ by $\bar{f}(\mathsf{M}_{G})=f(G)$. This is well-defined because $f$ is permutation-invariant and isolate-indifferent, as $\mathsf{M}_{G}=\mathsf{M}_{G^{\prime}}$ if and only if $G$ and $G^{\prime}$ differ by isolated vertices and node relabeling [28, Prop. 3]. Further, $\bar{f}$ is continuous by continuity of $f$, hence it extends uniquely to a continuous function on the closure $\mathcal{M}$ of $\mathcal{M}_{<\infty}$. This proves statement 2.
@@
For the first claim, it was shown in [31, Thm. 5.2] that there exists a polynomial $q_{k}$ over $\mathbb{R}^{k\times k}$ satisfying $f(G_{n})=\mathbb{E}q_{k}(\rho(F_{k,n})G_{n})$ for all $G_{n}\in\Delta^{n\times n}$ and all $n\in\mathbb{N}$. Since $q_{k}$ is $L$-Lipschitz on $\Delta^{k\times k}$ for some $L>0$, we have by Proposition 3.2 that
@@
We remark that the graphon analog of the characterization of quotient-testable parameters in Theorem 4.10 was used as the definition of testable parameters in [4, Def. 2.11].
@@
where $S_{2}^{(\mathrm{out})},S_{2}^{(\mathrm{in})}\in\mathbb{N}^{3\times 3}$ are outward- and inward-pointing stars with two leaves, respectively. This is an extension to weighted and directed graphs of the so-called first Zagreb index, originally introduced for chemical networks in [19]. Define the polynomial $q_{2}$ on $2\times 2$ matrices by
@@
We end this section by giving two further examples of quotient-testable parameters, each quantifying different structural properties.
Example 4.12 (Global Clustering Coefficient).
One extension of the global clustering coefficient to weighted undirected graphs, a global analog of the coefficient proposed in [48, §7] for gene co-expression networks, is
$f(G)=\frac{\mathrm{inj}(K_{3};G)}{\mathrm{inf}(P_{2};G)\max_{i,j}G_{i,j}},$ (16)
where $\mathrm{inj}(H;G)$ denotes the injective homomorphism number and is defined as in (2) but with the sum restricted to injective maps, $K_{3}$ is an arbitrarily-directed triangle graph, and $P_{2}$ is the directed path on three vertices. We proceed to show that (16) is quotient-testable. Proposition 4.9 shows that $\mathrm{inj}(K_{3};G)$ and $\mathrm{inj}(P_{2};G)$ extend to continuous functions on all of $\mathcal{M}$. Similarly, we note that $G\mapsto\max_{i,j}G_{i,j}$ extends to such a continuous function, since if $(G_{n})$ quotient-converges to a grapheur described by $(E,\sigma,\varsigma,\theta,\vartheta)$, then $\max_{i,j}(G_{n})_{i,j}\to\max_{i,j}E_{i,j}$ by Proposition 3.6. Finally, note that $f(G)\leq 1$ for all graphs $G$ of any size, proving that $f$ extends to a continuous function on all of $\mathcal{M}$ and hence is quotient-testable by Proposition 4.8.
Example 4.13 (Katz Centrality).
Another measure of importance, or influence, of a node, originally introduced in [30] for social networks, is its Katz centrality. Specifically, for a parameter $\alpha\in(0,1)$, we consider the sum of the Katz centralities of all the nodes
$f(G)=\mathbbm{1}^{\top}\Big[(I-\alpha G)^{-1}-I\Big]\mathbbm{1}=\sum_{k\geq 1}\alpha^{k}(\mathbbm{1}^{\top}G^{k}\mathbbm{1}),$
which is another parameter quantifying the presence of hubs in a network. Once again, this parameter is quotient-testable. Indeed, if $(G_{n})$ quotient-converges, then $\mathbbm{1}^{\top}G_{n}^{k}\mathbbm{1}$ converges for each $k$ by Proposition 4.9, hence $f(G_{n})$ converges as well for any fixed $\alpha\in(0,1)$.